OpenStudy (this):

-5k^2+10k-13=-8 a,0;two imaginary solution b,0;0ne real solution c,c,200;two real solution d,0;two real solution

OpenStudy (this):

some one help me plz

OpenStudy (tkhunny):

Is this just the discriminant, again? Add 8 and substitute away. Let's see what you get,

OpenStudy (tkhunny):

Did you add 8, yet? This wil make the right hand side zero (0). Zero is what you need. Nothing else will do.

OpenStudy (anonymous):

-5k^2+10k-13=-8 adding 8 on both sides gives -5k^2+10k-13+8=-8+8 -5k^2+10k-5=0 divide both sides by 5 to get -k^2+2k-1=0 multiply both sides by -1 gives k^2-2k+1=0 you can now factor this or use quadratic formula to find k

OpenStudy (anonymous):

k=?